OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the elimination method to solve the following system of equations. Part 2: Provide the solution to the system. 2x – 5y = –12 3x – 4y = 3

OpenStudy (anonymous):

PLEASE HELP!!! ASAP!!!! I'M SO LOST!

OpenStudy (anonymous):

you want to eliminate one of the variables from those two equations... are there two numbers you can think which, when each equation is multiplied by one of the numbers, allows you to eliminate a variable? eg. multiply the top equation by -3 and the bottom by ...?

OpenStudy (anonymous):

-3(2x -5y =-12) : -6x +15y = 36 +A( 3x -4y =3) what value of A would allow you to add the two equations and get rid of the 'x' terms?

OpenStudy (anonymous):

What? I'm confused..

OpenStudy (anonymous):

elimination means you're going to add the equations and eliminate one of the variables. just adding the two equations that you've been given won't do the trick, do you see that?

OpenStudy (anonymous):

so you scale each equation so that the coefficients on one of the variables will allow it to be eliminated when you add the equations. coefficient on 'x' in the top equation is now '-6' if the bottom equation had a coefficient on 'x' of '+6', you could add the equations and eliminate 'x'..

OpenStudy (anonymous):

-6x +15y = 36 +A( 3x -4y =3) -------------------- 0x so what should A be? (what should the bottom equation be multiplied by?)

OpenStudy (anonymous):

Nevermind my teacher just called & is explaining it, but thank you :)

OpenStudy (anonymous):

"I don't have a calculator & I don't know what you're talking about. I didn't learn this stuff very well. I have a rly bad teacher!"